OpenStudy (anonymous):

Find the volume of a sphere with a surface area of 1296π in^2

OpenStudy (anonymous):

Surface area of a sphere is 4pir^2. So, divide by 4pi: r^2 = 324, and r=18. Now plug that back in to volume formula, (4/3)*pi*r^3, and you'll get 7776pi.

OpenStudy (anonymous):

A=4πr^2 solve for r 1296π=4πr^2 324=r^2 r=18 V=4/3πr^3 V=4/3π(18^3) V=7776π in^3

OpenStudy (anonymous):

thanks you guys